OpenStudy (anonymous):

Identify the factors of 9x2 + 24x + 16. (3x + 4)(3x + 4) (3x − 4)(3x − 4) (3x + 2)(3x + 8) (3x − 8)(3x − 2)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Steps @Alice2200

OpenStudy (uri):

9x2 + 24x + 16. formula: (a+b)^2= a^2+2ab+b^2

OpenStudy (anonymous):

A; (3x + 4)(3x + 4) 3x * 3x = 9x^2 3x * 4 = 12x 4 * 3x = 12x 4 * 4 = 16. Add like terms: 9x^2 + 24x + 16

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

^^ Mikeyy1992 typed it faster.

OpenStudy (anonymous):

I'm sorry for my 136 words per minute speed.

OpenStudy (anonymous):

Yea, me too.

OpenStudy (anonymous):

Well, you have a nice day. Buh bye!(:

OpenStudy (anonymous):

Which polynomial is a difference of two squares?   x2 − 12 x2 + 12 x2 − 36 x2 + 36

OpenStudy (anonymous):

Can any of you help me with this one

OpenStudy (anonymous):

Let's see.

OpenStudy (anonymous):

First one is incorrect.

OpenStudy (anonymous):

Second one is incorrect; both don't have same squares.

OpenStudy (anonymous):

Third one is close, but is still incorrect. (x - 6)(x + 6) Last one is correct; (x +6)(x +6) <--- perfect square c:

OpenStudy (anonymous):

k Thanks

OpenStudy (anonymous):

Oh my bad, I mistaken your question.

OpenStudy (anonymous):

misread*

OpenStudy (anonymous):

The Third one is your answer. I thought question was asking for perfect square.

OpenStudy (anonymous):

Oh ok lol

OpenStudy (anonymous):

If you need help with another question, make a new one and tag me.